OpenStudy (anonymous):

Solve the following system. y = x + 3 3x + y = 19 A. (7, 4) B. (−4, 7) C. (4, 7) D. (4, −7)

OpenStudy (anonymous):

use substitution

OpenStudy (anonymous):

y = x + 3 3x + y = 19 -------------- y = ? so SUBSTITUTE WHAT " y = " in the second equation, like this: 3x + (x + 3) = 19 3x + x + 3 = 19 COMBINE LIKE TERMS (i.e. 3x + x) 3x + 1x + 3 = 19 4x + 3 = 19 SUBTRACT 3 FROM BOTH SIDES: 4x + 3 = 19 -3 -3 ---- ---- 4x = 16 DIVIDE BOTH SIDES BY 4 4x = 16 -- --- 4 4 x = ?

OpenStudy (e.cociuba):

y= x+3 you can plug the (y) in the other equation to get: ______________________ 3x +x+3 = 19 4x+3= 19 4x =19-3 4x = 16 x = 4 ________________________ Get (y) by plugging (x) into the equation: (see below) y = 4+3 y= 7